# Needed computations and computational capabilities for stellarators

### v1

@@ VIII.1 Possible divertor designs
The dictionary definition of chaos as complete disorder is misleading in the application of the concept to non-resonant divertors. Chaotic magnetic field lines can lie in well collimated flux tubes that extend from the plasma edge to the locations on the surrounding walls. These tubes define a non-resonant divertor, and the locations where they strike the walls are remarkably robust to changes in the plasma configuration Bader:2023 . Understanding how the location of the strike points and the width of the diverted flux tubes is controlled by the freedom in the external coils remains to be established.
@@
A strong bootstrap current is thought to make island divertors, the type of divertor used on W-7X, impractical. This may be less true in a power plant than in an experimental facility. A quasi-steady-state power plant could operate in a single plasma state with a known edge transform. Both the edge transform and surface shape can be controlled by external coils during startup. The rotational transform can be controlled by poloidal field coils forming a transformer that allows a controlled change in poloidal flux in the central hole of the torus.
VIII.2 A potentially advantageous stellarator configuration
@@
When the tritium density equals that of deuterium, the fraction of the tritium that is burned in an energy confinement time is a fraction of a percent. The tritium must make many passes through the plasma before it is burnt. This worsens a major issue for fusion power, the achievement of tritium self-sufficiency of deuterium-tritium power plants tritium ; tritium2 . Concepts for increasing the number of tritiums produced per neutron would have a major impact on the feasibility of fusion.
@@
Before the attractiveness of this concept can be assessed two physics issues must be better understood: (1) the physics of the interdiffusion of different species and (2) the localization and the depth of penetration of plasma deposition by pellets.
VIII.3 Other issues that should be studied
The consistency of various features of plasma configurations is important to study. An example is a large shear in a curl-free configuration with small neoclassical transport. Another is the tendency of optimized stellarators to have outer surfaces with high surface curvature. These regions are important for the formation of non-resonant divertors.
@@ VIII.3 Other issues that should be studied
A related issue that should be studied is whether stellarators can be designed so Alfvénic modes can be driven by the energetic alpha particles but only in the central region. Central Alfvénic modes do not directly cause an enhancement of the energetic alpha losses to the walls. To the extent these alpha-driven waves are damped by the viscosity associated with background deuterium and tritium ions, direct ion heating is achieved. Normally, the alpha energy is transferred to the electrons, which indirectly transfer it to the ions. This indirect method fails to heat ions when the electron-ion equilibration time is long compared to the ion energy confinement time, which can be the case when the electrons become too hot Boozer:omnigeneity .
IX Effects of sudden plasma termination
@@ Appendix A Freedom in external magnetic fields
Two time-independent equations, $\vec{\nabla}\cdot\vec{B}=0$ and ×$\vec{\nabla}\times\vec{B}=\mu_{0}\vec{j}$, define the full freedom of magnetic fields that can be produced in a plasma chamber by external coils. Within the region enclosed by a toroidal surface that lies closer to the plasma than any coil, the coil-produced magnetic field must have the form $\vec{B}=\vec{\nabla}\phi_{\ell}$, where $\phi_{\ell}$ is a solution to Laplace’s equation, $\nabla^{2}\phi_{\ell}=0$.
The general solution $\phi_{\ell}$ in a toroidal region is the sum of a single-valued term $\phi$ and two multivalued terms. One multivalued term increases by $\mu_{0}G_{0}$ each toroidal circuit of the torus and the other by $\mu_{0}I_{0}$ each poloidal circuit. A non-zero $G_{0}$, which is the net current coming through the hole in the center of the toroidal enclosing surface, determines the toroidal flux within that surface. A non-zero $I_{0}$, which is the net toroidal current outside the enclosing surface, determines the magnetic flux through the hole in the center of the enclosing surface. The effect of $I_{0}$ within the enclosing surface can represented by the single-valued potential $\phi$.
@@ A.3 Specification of external magnetic field
The external magnetic field has two roles: (1) The field from the basic coil set $\vec{B}_{p}$ must provide the toroidal magnetic flux within the plasma. For that, at least one coil must encircle the plasma poloidally. (2) The dipole field must supply whatever normal magnetic field is required on the desired plasma surface to ensure that surface is a magnetic surface. Other than providing the toroidal flux, the coils producing $\vec{B}_{p}$ are unconstrained.
@@
When the entire bounding surface is covered with the grid of dipoles, then a solution will be obtained with the smallest singular value as large as possible with a given coil-plasma separation. A discussed by Todd Elder Elder:thesis2023 ; Elder-Boozer:2024 , some fraction $f_{d}$ of the dipoles can be eliminated from the $\tensor{M}$, the ones with the smallest magnitudes or the ones most inconveniently located, and a new SVD calculated. As long as the smallest singular value remains sufficiently large, the remaining patches of dipoles can accurately produce the normal magnetic field on the plasma surface to the required accuracy.
@@
The fineness of the required dipole grid is largely determined by the number of magnetic field distributions are associated with sufficiently large singular values to be of practical importance. The magnetic field itself is accurately represented by a relatively course grid. The grid is defined by the length $\delta_{\ell}$ of a side of a grid cell $\ell$. Following a 2011 discussion by Boozer and Ku Boozer-Ku:control2011 , the approximate answer is $\delta_{\ell}\lesssim\Delta^{(p)}_{\ell}/1.3$, where $\Delta^{(p)}_{\ell}$ is the distance cell $\ell$ is from the plasma surface, on which the external magnetic field is to be controlled. A finer grid may be needed to calculate the required surface currents on the bounding surface with the desired accuracy
@@ Acknowledgements
This material is based upon work supported by the grant 601958 within the Simons Foundation collaboration Hidden Symmetries and Fusion Energy and by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences under Awards DE-FG02-95ER54333 and DE-FG02-03ER54696.

### v2

@@ IV.1 Possible divertor designs
The dictionary definition of chaos as complete disorder is misleading in the application of the concept to non-resonant divertors. Chaotic magnetic field lines can lie in well collimated flux tubes that extend from the plasma edge to the locations on the surrounding walls. These tubes define a non-resonant divertor, and the locations where they strike the walls are remarkably robust to changes in the plasma configuration Bader:2023 .
Non-resonant divertors appear naturally in optimized stellarators due to locations of sharp curvature of the magnetic surfaces near the plasma edge, Bader et al Bader:2018 . Due to $\vec{\nabla}\cdot\vec{B}=0$, these are also the locations at which the separation between neighboring magnetic surfaces becomes large. Boozer and Punjabi Boozer-Punjabi:2018 used these features to derive an analytic magnetic-field-line Hamiltonian, the poloidal magnetic flux as a function to the toroidal magnetic flux and two angles, $\psi_{p}(\psi,\theta,\varphi)$. They have used this Hamiltonian in studies of topological features of non-resonant divertors, such as Cantori, turnstiles, and magnetic field line chaos.
@@
The magnetic surfaces of the well-confined plasma are separated from the sea of chaotic field lines that surrounds it by Cantori, which are like irrational magnetic surfaces that have localized holes through which magnetic field lines can pass, with equal fluxes in both directions, and reach the chamber walls. These holes are caused by nearby unstable fixed points—field lines that exactly close on themselves after making a number of transits of the torus but from which neighboring lines exponentiate. In helical symmetry, these would be the X-points of islands, but in the absence of symmetry, the field lines in the neighborhood of an X-point become chaotic.
Using a map that models a non-resonant divertor, Boozer and Punjabi Boozer-Punjabi:2018 found that the probability that a field line will be lost from a Cantorus in one toroidal transit has the form
$P_{\ell}(\psi)=\left(\frac{\psi-\psi_{o}}{w_{\ell}}\right)^{d},$ (4)
where $\psi$ is the toroidal flux enclosed by the Cantorus, $\psi_{o}$ is the flux enclosed by the outermost magnetic surface, $w_{\ell}$ is a constant, and $d\approx 2$. But, the outward and inward flux tubes through a Cantorus need not be adjacent. Adjacent versus non adjacent turnstiles can have different $d$’s Punjabi:2022 .
The number of toroidal transits required for a field line to be lost can be written as
$N_{\ell}(\psi)=\left(\frac{w_{\ell}}{\psi-\psi_{o}}\right)^{d}.$ (5)
The number of toroidal transits that would be required for plasma to diffuse to the Cantorus at which a field line is lost is $N_{\ell}(\psi_{\ell})=(\psi_{\ell}-\psi_{o})^{2}/D_{p}$. Equating the $N_{\ell}$ from this equation with that from Equation (5) gives
$\psi_{\ell}-\psi_{o}$ $=$ $\left(D_{p}w_{\ell}^{d}\right)^{\frac{1}{2+d}};$ (6)
$N_{\ell}$ $=$ $\left(\frac{w_{\ell}^{2}}{D_{p}}\right)^{\frac{d}{2+d}};$ (7)
$P_{\ell}$ $=$ $\frac{1}{N_{\ell}}.$ (8)
$P_{\ell}$ gives the size of each of the two escaping flux tubes that carries plasma to the walls. Each has a magnetic flux is $\delta\psi=\psi_{\ell}/N_{\ell}$.
In a non-resonant divertor the plasma flows to the wall along one or more pairs of flux tubes, which appear to be very narrow. The compactness of the flux tubes that carry the plasma to the divertor may seem surprising. The same narrowness is important in the runaway electron problem in post-disruption tokamaks. Relativistic electrons from a central region of chaotic field lines exit the plasma along narrow flux tubes when the field evolves slowly compared to the time required for a relativistic electron to cover the chaotic region but exits the plasma over a broad region when the evolution is fast runaway:2016 . Both effects are seen in tokamak experiments.
A way to study the collimation of the narrow flux tubes carrying plasma to the divertor is to cover the surrounding chamber wall with a grid. In almost all grid cells, the normal magnetic field to the wall will be non-zero, and the field lines can be followed from the location where they enter the chamber to the place where they exit after $N_{t}$ toroidal transits. Equation (8) says the probability of a field line staying in as $N_{t}\rightarrow\infty$ is zero. Nonetheless there can be pairs of points on the chamber wall where field lines make $N_{t}\rightarrow\infty$ transits. These lines with $N_{t}\rightarrow\infty$ come extremely close to a magnetic surface, which can be the outermost magnetic surface of the confining plasma and are the lines to be used in divertors. Another possibility is that lines with $N_{t}\rightarrow\infty$ come extremely close to the outermost surface of a magnetic island that lies outside the plasma.
In a stellarator the helical field $B_{h}$ is much stronger than the net poloidal field $B_{\iota}\propto(B_{h}/B)^{2}B$ that determines the rotational transform Spitzer:1958 . In an unpublished study, Alkesh Punjabi finds that in most grid cells field lines quickly exit, $N_{t}\sim 1$. Although the magnetic field lines are in principle chaotic in the region outside those bounded by magnetic surfaces, once a magnetic field line escapes from the Cantori that surround these surfaces, the line strikes the wall within a few toroidal transits, so the exponential separation of neighboring lines does not occur over a sufficient length to destroy the compactness of the exiting flux tubes.
Understanding how the location of the strike points and the width of the diverted flux tubes are controlled by the freedom in the external coils remains to be established. Studies of curl-free stellarator configurations would clarify how the location of the $N_{t}\rightarrow\infty$ points and the widths the holes in the Cantori, the $w_{\ell}$, are determined. The control of both is very important for the design of attractive non-resonant divertors. Although the location of the $N_{t}\rightarrow\infty$ points is remakably robust, baffles to keep neutrals close to divertor pumps can be better designed the more accurately the location of these points can be defined and the behavior of field lines near these points can be determined, which is related to $w_{\ell}$.
IV.2 A potentially advantageous stellarator configuration
@@
When the tritium density equals that of deuterium, the fraction of the tritium that is burnt in an energy confinement time is a fraction of a percent. This follows from the enormous ratio of the energy of the fusion-produced alpha particles to the temperature of the plasma. The tritium must make many passes through the plasma before it is burnt. This worsens a major issue for fusion power, the achievement of tritium self-sufficiency of deuterium-tritium power plants tritium ; tritium2 . Concepts for increasing the number of tritiums produced per neutron would have a major impact on the feasibility of fusion.
@@
In standard stellarator reactor designs, the helium and high-Z impurities are flushed out of the plasma by a 50-50 deuterium-tritium mixture. Tritium is many orders of magnitude more expensive than deuterium. The zoned concept tries to avoid the bizarre use of tritium as a flushing fluid.
Before the attractiveness of the zoned concept can be assessed two physics issues must be better understood: (1) the physics of the interdiffusion of different species and (2) the localization and the depth of plasma penetration by pellets.
IV.3 Other issues that should be studied
The consistency of various features of plasma configurations is important to study. An example, is it possible to have a large shear in the rotational transform in a curl-free configuration with small neoclassical transport? Another is the tendency of optimized stellarators to have outer surfaces with high surface curvature. These regions are important for the formation of non-resonant divertors.
@@ IV.3 Other issues that should be studied
A related issue that should be studied is whether stellarators can be designed so Alfvénic modes can be driven by the energetic alpha particles but only in the central region. Central Alfvénic modes do not directly cause an enhancement of the energetic alpha losses to the walls. To the extent these alpha-driven Alfvén waves are damped by the viscosity associated with background deuterium and tritium ions, direct ion heating is achieved. Normally, the alpha energy is transferred to the electrons, which indirectly transfer it to the ions. This indirect method fails to heat ions when the electron-ion equilibration time is long compared to the ion energy confinement time, which can be the case when the electrons become too hot Boozer:omnigeneity .
V Effects of sudden plasma termination
@@ Appendix A Freedom in external magnetic fields
Two time-independent equations, $\vec{\nabla}\cdot\vec{B}=0$ and ×$\vec{\nabla}\times\vec{B}=\mu_{0}\vec{j}$, define the full freedom of magnetic fields that can be produced in a plasma chamber by external coils. Within the region enclosed by a toroidal surface that lies closer to the plasma than any coil, the coil-produced external magnetic field must have the form $\vec{B}_{x}=\vec{\nabla}\phi_{\ell}$, where $\phi_{\ell}$ is a solution to Laplace’s equation, $\nabla^{2}\phi_{\ell}=0$.
The general solution $\phi_{\ell}$ in a toroidal region is the sum of a single-valued term $\phi$ and two multivalued terms, which depend linearly on $\theta$ or $\varphi$. One multivalued term increases by $\mu_{0}G_{0}$ each toroidal circuit of the torus and the other by $\mu_{0}I_{0}$ each poloidal circuit. A non-zero $G_{0}$, which is the net current coming through the hole in the center of the toroidal enclosing surface, determines the toroidal flux within that surface. A non-zero $I_{0}$, which is the net toroidal current outside the enclosing surface, determines the magnetic flux through the hole in the center of the enclosing surface. The effect of $I_{0}$ within the enclosing surface can represented by the single-valued potential $\phi$.
@@ A.3 Specification of external magnetic field
The external magnetic field has two roles: (1) The field from the basic coil set $\vec{B}_{p}$ must provide the toroidal magnetic flux. For that, at least one coil must encircle the plasma poloidally. (2) The dipole field must supply whatever normal magnetic field is required on the desired plasma surface to ensure that surface is a magnetic surface. Other than providing the toroidal flux, the coils producing $\vec{B}_{p}$ are unconstrained.
@@
When the entire bounding surface is covered with the grid of dipoles, then a solution will be obtained with the smallest singular value as large as possible with a given coil-plasma separation. As discussed by Todd Elder Elder-Boozer:2024 , some fraction $f_{d}$ of the dipoles can be eliminated from the $\tensor{M}$, the ones with the smallest magnitudes or the ones most inconveniently located, and a new SVD calculated. As long as the smallest singular value remains sufficiently large, the remaining patches of dipoles can accurately produce the normal magnetic field on the plasma surface to the required accuracy.
@@
The fineness of the required dipole grid is largely determined by the number of magnetic field distributions are associated with sufficiently large singular values to be of practical importance. The magnetic field itself is accurately represented by a relatively coarse grid. The grid is defined by the length $\delta_{\ell}$ of a side of a grid cell $\ell$. Following a 2011 discussion by Boozer and Ku Boozer-Ku:control2011 , the approximate answer is $\delta_{\ell}\lesssim\Delta^{(p)}_{\ell}/1.3$, where $\Delta^{(p)}_{\ell}$ is the distance cell $\ell$ is from the plasma surface, on which the external magnetic field is to be controlled. A finer grid may be needed to calculate the required surface currents on the bounding surface with the desired accuracy
@@ Acknowledgements
This material is based upon work supported by the grant 601958 within the Simons Foundation collaboration Hidden Symmetries and Fusion Energy and by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences under Awards DE-FG02-95ER54333 and DE-FG02-03ER54696 and Sub-Awards Princeton University PU SUB0000764 and Princeton Plasma Physics Laboratory PPPL P-240000750.